OpenStudy (anonymous):

What is the absolute value of the number? 3 + i *square root* 10 4 8 *square root* 3 ( i think its the last , but im not sure ? )

OpenStudy (anonymous):

Is he number a negative or positive?

OpenStudy (helder_edwin):

if \(z=a+bi\) then \[ \large |z|=\sqrt{a^2+b^2} \]

OpenStudy (anonymous):

the*

OpenStudy (anonymous):

@andreascott you thought correct

OpenStudy (anonymous):

thanks @soty2013

OpenStudy (helder_edwin):

it is not the fourth choice.

OpenStudy (helder_edwin):

use the definition i posted